OpenStudy (anonymous):

PLEASE HELP ME. I NEED YOU TO EXPLAIN Find the curvature of r(t)= at the point (1,0,0). How do I know which of that points I need to use? The correct answer must be t=1 but why???

OpenStudy (anonymous):

The curvature \(\kappa\) of a vector function \(\vec{r}(t)\) is given by \[\kappa=\frac{\left\|\vec{r}'(t)\times\vec{r}''(t)\right\|}{\left\|\vec{r}'(t)\right\|}\] You have \[\begin{cases} \vec{r}(t)=\left\langle t^2,\ln t,t\ln t\right\rangle\\\\ \vec{r}'(t)=\left\langle 2t,\dfrac{1}{t},\ln t+1\right\rangle\\\\ \vec{r}(t)=\left\langle 2,-\dfrac{1}{t^2},\dfrac{1}{t}\right\rangle \end{cases}\] with \[\begin{align*}\vec{r}'(t)\times\vec{r}''(t)&=\begin{vmatrix}\vec{i}&\vec{j}&\vec{k}\\ 2t&\frac{1}{t}&\ln t+1\\ 2&-\frac{1}{t^2}&\frac{1}{t}\end{vmatrix}\\\\ &=\begin{vmatrix}\frac{1}{t}&\ln t+1\\-\frac{1}{t^2}&\frac{1}{t}\end{vmatrix}\vec{i}-\begin{vmatrix}2t&\ln t+1\\ 2&\frac{1}{t}\end{vmatrix}\vec{j}+\begin{vmatrix}2t&\frac{1}{t}\\ 2&-\frac{1}{t^2}\end{vmatrix}\vec{k}\\\\ &=\left\langle\frac{2+\ln t}{t^2},2\ln t,-\frac{4}{t}\right\rangle \end{align*}\] and so on.

OpenStudy (anonymous):

As for which value of \(t\) to use: you're asked to find the curvature at the point \((1,0,0)\), which you can use to find \(t\). What value of \(t\) gives \[\begin{cases} t^2=1\\ \ln t=0\\ t\ln t=0 \end{cases}~~?\]